# IndoCulture: Exploring Geographically-Influenced Cultural Commonsense Reasoning Across Eleven Indonesian Provinces

### v1

@@ 5 Discussion
The recent study Wang et al. (2024) has demonstrated that evaluating language models using token probability in multiple-choice question types does not align well with the generated text. This discrepancy is reported to be more pronounced in models fine-tuned on conversational or safety data. In response to this issue, we have conducted a manual evaluation on 200 random samples (in Section 4.3) and found that the performance in generated texts, especially for Merak, the best open-source language model, is deteriorating. However, this issue is less apparent in GPT-4, as the initial evaluation is based on regular expressions. Conducting manual evaluation on all data and models is expensive, and we plan to address this issue in future works. This work primarily focuses on introducing a novel dataset meticulously constructed for evaluating cultural commonsense reasoning within the Indonesian context, including preliminary evaluation results based on the standard methods used in previous studies OpenAI (2023); Touvron et al. (2023); Koto et al. (2023); Li et al. (2023b); Koto et al. (2024).
@@ 6 Conclusion
IndoCulture is a cultural commonsense reasoning dataset encompassing the diversity of Indonesian cultures, spanning from Aceh province in the west to Papua province in the east. Through collaboration with local individuals across eleven provinces and rigorous quality control measures, we introduce IndoCulture for the purpose of evaluating language models. Our findings reveal that large language models, whether Indonesian-centric or multilingual, demonstrate a limited understanding of Indonesian culture. Notably, incorporating location as additional context significantly enhances model performance, particularly in GPT–3.5 and GPT–4.
@@
IndoCulture is specifically designed to explore the influence of geographical location on cultural commonsense reasoning, with a focus on the present time. It does not consider temporal aspects. Our dataset was created in the year 2023, recognizing that cultures may evolve over time, as discussed by Mesoudi (2016).
@@
IndoCulture is licensed under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License. [^10] https://creativecommons.org/licenses/by-nc-sa/4.0/ Our data is intended for academic research and non-commercial purposes. Workers are compensated above the minimum monthly salary in Indonesia and are fully aware that the data will be released to the public. It is important to note that no private or sensitive information of the workers is included in IndoCulture.
Appendix A Model Artifacts
@@
LLamA2-chat (7B) meta-llama/Llama-2-7b-chat
LLamA2 (13B) meta-llama/Llama-2-13b
LLamA2-chat (13B) meta-llama/Llama-2-13b-chat
@@
Table 6: With the exception of GPT-3.5 and GPT-4, all the models used in this study were sourced from Huggingface Wolf et al. (2020).

### v2

@@ 5 Discussion
A recent study Wang et al. (2024) has demonstrated that evaluating language models using token probabilities in multiple-choice question types does not align well with the generated text. This discrepancy is reported to be more pronounced in models fine-tuned on conversational or safety data. In response to this issue, we conducted a manual evaluation of 200 random samples (in Section 4.3) and found that the performance in the generated text, especially for Merak — the best open-weight Indonesian-centric language model — deteriorates. However, this issue is less apparent in larger models such as Llama–3 and GPT–4. Conducting manual evaluation on all data and models is expensive, and we plan to address this issue in future works. This work primarily focuses on introducing a novel dataset constructed for evaluating cultural commonsense reasoning within the Indonesian context, including preliminary evaluation results based on standard methods used in previous studies OpenAI (2023); Touvron et al. (2023); Koto et al. (2023); Li et al. (2024); Koto et al. (2024).
@@ 6 Conclusion
IndoCulture is a cultural commonsense reasoning dataset encompassing the diversity of Indonesian cultures, spanning from Aceh province in the west to Papua province in the east. Through collaboration with local individuals across eleven provinces and rigorous quality control measures, we introduce IndoCulture for the purpose of evaluating language models. Our findings reveal that large language models, whether Indonesian-centric or multilingual, demonstrate a limited understanding of Indonesian cultures. Notably, incorporating location as additional context significantly enhances model performance, particularly for GPT–3.5 and GPT–4.
@@
IndoCulture is specifically designed to explore the influence of geographical location on cultural commonsense reasoning, with a focus on the present time. It does not consider temporal aspects. Our dataset was created in the year 2023, and we recognize that cultures may evolve over time, as discussed by Mesoudi (2016).
@@
We also acknowledge that our dataset size is relatively small, to cover all the provinces. However, compared to existing Indonesian datasets that have been manually curated by natives (see COPAL–ID in Table 1), ours is significantly larger in terms of both size and regional coverage. Future work may extend the size and the coverage of IndoCulture to get a more holistic picture of Indonesian cultures.
@@
IndoCulture is licensed under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License. [^8] https://creativecommons.org/licenses/by-nc-sa/4.0/ Our data is intended for academic research and non-commercial purposes. Workers were compensated above the minimum monthly salary in Indonesia and are fully aware that the data will be released to the public. It is important to note that no private or sensitive information of the workers is included in IndoCulture.
Acknowledgements
We acknowledge the thorough feedback and impactful suggestions from the reviewers and the action editor. We also extend our gratitude to the 22 annotators from eleven provinces in Indonesia for their valuable contributions in constructing IndoCulture. This project was supported by MBZUAI and UI through PUTI-Q2 grant (NKB-1192/UN2.RST/HKP.05.00/2022).
Appendix A Model Details
@@
Llama–2 chat (7B) meta-llama/Llama--2-7b-chat
Llama–2 (13B) meta-llama/Llama--2-13b
Llama–2 chat (13B) meta-llama/Llama--2-13b-chat
[…]
Llama–3 Instruct (8B) meta-llama/Meta-Llama--3-8B-Instruct
Llama–3 (70B) meta-llama/Meta-Llama--3-70B
Llama–3-chat (70B) meta-llama/Meta-Llama--3-70B-Instruct
@@
Table 7: With the exception of GPT–3.5 and GPT–4, all the models used in this study were sourced from Huggingface Wolf et al. (2020).